# Semantic Density: Uncertainty Quantification in Semantic Space for Large Language Models

### v2

@@ 3.4 Dimension-invariant Kernel
In the semantic density estimator use case, one drawback of the original Epanechnikov kernel is that the normalization coefficient $\frac{\Gamma(2+\frac{D}{2})}{\pi^{\frac{D}{2}}}$ changes with the dimension $D$ of $\bm{v}$. As a result, semantic densities calculated using embeddings with different dimensionalities are incomparable. This issue may limit the flexibility in selecting embedding methodologies for semantic density calculation. However, the normalization coefficient can be removed to make the kernel function simpler and more flexible, without affecting the performance of uncertainty quantification. The kernel function of Eq. 6 thus becomes
@@
Although the resulting kernel function does not meet the normalization requirement in standard KDE, it fits uncertainty quantification well. As long as the norm requirements in Eq. 2,  3 and  4 are fulfilled, any embedding models can be used to generate $\bm{v}$, regardless of the embedding dimensionalities. The outcome of kernel function is always within $[0,1]$, and a kernel value of 1, $\frac{1}{2}$, and 0 correspond to semantically equivalent, irrelevant and contradictory responses, respectively. As a result, the semantic density is also within $[0,1]$, with 1 as the highest semantic density a response can obtain, indicating that all the reference responses are semantically equivalent to it; analogously, it obtains a semantic density of 0 when all reference responses are semantically contradictory to it. This consistency in the value range makes practical applications of semantic density convenient: Practitioners can set a fixed threshold on semantic density to detect unreliable responses.
@@ 3.5 Semantic Distance Measurement via the Natural Language Inference (NLI) Model
Although most of the existing text-embedding models work in the semantic space defined in Section 3.2, they do not perform well in measuring semantic similarities [26]. Moreover, they can only consider input texts as a whole instead of doing a contextual inference on part of the input, i.e., they can only obtain $E(\bm{x}+\bm{y})$ instead of $E(\bm{y}|\bm{x})$, where $\bm{x}+\bm{y}$ means a concatenation of $\bm{x}$ and $\bm{y}$.
The natural language inference (NLI) classification model[9] has proven to be effective in analyzing the semantic relationship between LLM responses with the prompt as context [16]. Given a pair of texts, a NLI model performs a classification task and outputs the probabilities for them to be semantically equivalent ("entailment" class), irrelevant ("neutral" class), or contradictory ("contradiction" class). Given the output class probabilities, the expectation of $||\bm{v}_{*}-\bm{v}_{i}||$ can be obtained as
@@
where $p_{\mathrm{c}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$, $p_{\mathrm{n}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ and $p_{\mathrm{e}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ are the probabilities for $\bm{y}_{*}$ and $\bm{y}_{i}$ to be semantically contradictory ("c" for "contradiction" class), irrelevant ("n" for "neutral" class) and equivalent ("e" for "entailment" class), respectively, given context $\bm{x}$. The expected value of $||\bm{v}_{*}-\bm{v}_{i}||^{2}$ can then be used in Eq. 8 to obtain the kernel function output.
@@ 3.6 Summary of the Semantic Density Framework
0:   $\bm{y}_{*}$: target response that needs uncertainty estimation $\bm{x}$:   original prompt for generating $\bm{y}_{*}$ $M$: number of unique reference responses to be sampled given $\bm{x}$
@@ 4 Experiments
This section first evaluates the performance of semantic density by comparing it with six existing uncertainty metrics over various LLMs and benchmarks. After that, two empirical studies are performed to investigate the robustness of semantic density when the number of reference responses and sampling strategy for target response are varied.
@@ 4.1 Performance Evaluation
Table 1: Performance of different uncertainty metrics across various LLMs and datasets
@@
Following the usual evaluation approach in the literature [16], the uncertainty metric is used as a quantitative indicator of how likely the response is going to be correct. Uncertainty values above a threshold are taken as incorrect while those below are taken as correct. For each threshold, the true positive rate vs. false positive rate is then measured. The area under this curve, namely area under receiver operator characteristic curve (AUROC), is calculated for each uncertainty metric. The AUROC score equals the probability that a randomly chosen incorrect response has a higher uncertainty than a randomly chosen correct response. A perfect uncertainty metric would have an AUROC score of 1 while a random metric would have 0.5. Note also that a higher semantic density corresponds to a lower uncertainty.
@@
The performance of semantic density (SD) was compared with six existing LLM uncertainty quantification methods: semantic entropy (SE) [16], P(True) [15], degree (Deg) [20], length-normalized likelihood (NL) [25], length-normalized entropy (NE) [23] and predictive entropy (PE) [15]. These methods were applied to seven SOTA open-source LLMs: Llama-2-13B [37], Llama-2-70B [37], Llama-3-8B [2], Llama-3-70B [2], Mistral-7B [12], Mixtral-8x7B [13] and Mixtral-8x22B [35]. Each LLM was tested on four free-form question-answering datasets commonly used in the literature: CoQA [30], TriviaQA [14], SciQ [40] and Natural Questions (NQ) [17]. For each question, 10 responses were generated using group beam search and used as reference responses in calculating SD, SE, Deg, NE and PE (note that P(True) and NL do not need reference responses). Each unique response among these 10 will also be used as a target response, e.g., the response that needs an uncertainty estimation, in calculating the AUROC scores of uncertainty metrics. Detailed experimental configuration and parametric setup is described in Appendix A.1.
@@
Table 1 shows the AUROC scores of each uncertainty metric across different models and datasets, with the best entry in each configuration highlighted in boldface. SD performs best in 26 out of 28 cases, demonstrating that it is reliable and robust as an uncertainty metric for LLM responses. In two cases it is outperformed by Deg. After investigation, the inherent sequence likelihood returned by the original LLM was badly calibrated in these two cases. Deg is the only method that ignores the likelihood information during its calculation, making its performance unaffected by this negative factor. However, for the other 26 cases, SD is able to utilize the likelihood information to its advantage, and outperform Deg.
@@
To confirm that the observed differences are statistically significant, a paired $t$-test (paired by LLM and dataset) was performed between SD and the other metrics (Table 2 in Appendix A.2). The $p$-values are consistently below $10^{-6}$, indicating that the performance gains of SD are strongly statistically significant.
@@ 4.2 Robustness of Semantic Density
In real-world applications, users may have different preferences when using LLMs to generate responses: Some may prefer a greedy sampling strategy while others may need diverse responses. The second study thus investigated how each uncertainty metric performs when the target response is sampled using different such strategies. The diverse beam search method inherently utilizes different strategies for each beam group: The first group performs a greedy beam search while later groups encourage more diverse responses. Following the setup in Section 4.1, the AUROC scores were calculated for target responses from each group separately, and the results averaged over the four datasets.
@@
Figure 2: AUROC scores over the different beam groups. The plots show the average AUROC scores over the four datasets for the different beam groups in diverse beam search. Smaller group index corresponds to the group with more greedy generation strategy, while the group with larger index tends to be more diverse during response generation. Each subfigure corresponds to one of the seven LLMs, as indicated by the subfigure heading. Each curve represents one uncertainty metric indicated by the legend at lower right. Semantic density exhibits consistently better and stable performance across different groups, compared to other methods.
@@ 5 Discussion and Future work
One limitation of semantic density is that it needs access to the output probabilities of generated tokens, which may not be available in some proprietary LLMs. In such a case, the more expensive variant in Eq. 5 can be considered as an alternative. Since semantic density does not require any further access to the internal states or weights of the original LLMs, it is still widely applicable.
The framework for measuring semantic density is modular, and therefore its main components can be extended in future work. First, new strategies that explicitly encourage a better coverage of semantic space can be developed to generate reference responses. Such extensions should improve the reliability of semantic density further. Second, text embedding methods that can measure contextual semantic similarity between responses more reliably will be helpful as well. Third, kernel functions specifically designed for the semantic space should allow utilizing semantic relationships more efficiently. Fourth, more precise methods for calibrating inherent token probabilities will form a more reliable base for calculating semantic density.
@@ 6 Conclusion
This paper proposes semantic density as a practical new metric for measuring uncertainty in LLM responses. It overcomes the limitations of existing approaches by utilizing the semantic information in an efficient and precise manner. It is response-specific, "off-the-shelf", and applicable to free-form generation tasks. Experimental comparisons with six existing uncertainty metrics across seven SOTA LLMs and four benchmark datasets suggest that it is accurate, robust, and general, and can therefore help deploy LLMs in safety-critical domains.
@@ A.1 Experimental Setup
This section provides the detailes of experimental setup for reproducing the results presented in Section 4.
@@
Following the setup in Kuhn et al. [16], for all the datasets, an answer is considered to be correct if its Rouge-L [18] to any of the reference answers is larger than 0.3, after trimming the redundant continuations. The details of the datasets are as follows:
@@
Uncertainty Metrics:
The exactly same target response and reference responses are used for all the tested methods. For SD, SE and Deg, the microsoft/deberta-large-mnli model from Huggingface Transformers library [41] is used as the NLI classification model. Following Kuhn et al. [16], the probabilities for "contradiction", "neutral" and "entailment" are averaged bidirectionally. The parametric setup for the uncertainty metrics tested in Section 4 is summarized as follows:
  • semantic density (SD): The exactly same steps as described by Algorithm 1 were implemented, with a fixed temperature of 0.1 applied to rescale the value of each token probability during postprocessing.
@@
  • degree (Deg): The "entailment" probability returned by the NLI model (averaged bidirectionally) is used as the similarity between two responses, which is then used as the diagonal element in degree matrix.
@@ A.2 Results of Statistical Tests
Table 2 shows the $p$-values of the paired $t$-tests as described in Section 4.1.
Table 2: Statistical significance of SD’s advantage over other methods ($p$-values of paired $t$-tests)

### v3

@@ 3.4 Dimension-invariant Kernel
In the semantic density estimator use case, one drawback of the original Epanechnikov kernel is that the normalization coefficient $\frac{\Gamma(2+\frac{D}{2})}{\pi^{\frac{D}{2}}}$ changes with the dimension $D$ of $\bm{v}$. As a result, semantic densities calculated using embeddings with different dimensionalities are incomparable. This issue may limit the flexibility in selecting embedding methodologies for semantic density calculation. However, the normalization coefficient can be removed to make the kernel function simpler and more flexible, without affecting the performance of confidence measurement. The kernel function of Eq. 6 thus becomes
@@
Although the resulting kernel function does not meet the normalization requirement in standard KDE, it fits confidence estimation well. As long as the norm requirements in Eq. 2,  3 and  4 are fulfilled, any embedding models can be used to generate $\bm{v}$, regardless of the embedding dimensionalities. The outcome of the kernel function is always within $[0,1]$, and a kernel value of 1, $\frac{1}{2}$, and 0 correspond to semantically equivalent, irrelevant, and contradictory responses, respectively. As a result, the semantic density is also within $[0,1]$, with 1 as the highest semantic density a response can obtain, indicating that all the reference responses are semantically equivalent to it; analogously, it obtains a semantic density of 0 when all reference responses are semantically contradictory to it. This consistency in the value range makes practical applications of semantic density convenient: Practitioners can set a fixed threshold on semantic density to detect unreliable responses.
@@ 3.5 Semantic Distance Measurement via the Natural Language Inference (NLI) Model
Although most of the existing text-embedding models work in the semantic space defined in Section 3.2, they do not perform well in measuring semantic similarities [33]. Moreover, they can only consider input texts as a whole instead of doing a contextual encoding on part of the input, i.e., they can only obtain $E(\bm{x}+\bm{y})$ instead of $E(\bm{y}|\bm{x})$, where $\bm{x}+\bm{y}$ means a concatenation of $\bm{x}$ and $\bm{y}$.
The natural language inference (NLI) classification model[15] has proven to be effective in analyzing the semantic relationship between LLM responses with the prompt as context [23]. Given a pair of texts, an NLI model performs a classification task and outputs the probabilities for them to be semantically equivalent ("entailment" class), irrelevant ("neutral" class), or contradictory ("contradiction" class). Given the output class probabilities, the expectation of $||\bm{v}_{*}-\bm{v}_{i}||$ can be obtained as
@@
where $p_{\mathrm{c}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$, $p_{\mathrm{n}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ and $p_{\mathrm{e}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ are the probabilities for $\bm{y}_{*}$ and $\bm{y}_{i}$ to be semantically contradictory ("c" for "contradiction" class), irrelevant ("n" for "neutral" class) and equivalent ("e" for "entailment" class), respectively, given context $\bm{x}$. During implementation, each response $\bm{y}$ will be concatenated with its prompt $\bm{x}$ (with prompt placed before the response) to form one text, i.e., $\bm{x}+\bm{y}$. Each input of the NLI model will then be a pair of these texts, analyzing the semantic relationship between two responses given the prompt. The expected value of $||\bm{v}_{*}-\bm{v}_{i}||^{2}$ can then be used in Eq. 8 to obtain the kernel function output.
@@ 3.6 Summary of the Semantic Density Framework
0:   $\bm{y}_{*}$: target response that needs confidence measurement $\bm{x}$:   original prompt for generating $\bm{y}_{*}$ $M$: number of unique reference responses to be sampled given $\bm{x}$
@@ 4 Experiments
This section first evaluates the performance of semantic density by comparing it with six existing uncertainty/confidence metrics over various LLMs and benchmarks. After that, two empirical studies are performed to investigate the robustness of semantic density when the number of reference responses and sampling strategy for target response are varied.
@@ 4.1 Performance Evaluation
Table 1: Performance of different uncertainty/confidence metrics across various LLMs and datasets
@@
Following the usual evaluation approach in the literature [23], the uncertainty/confidence metric is used as a quantitative indicator of how likely the response is going to be correct. Uncertainty values above a threshold are taken as incorrect while those below are taken as correct (and vice versa for confidence scores). For each threshold, the true positive rate vs. false positive rate is then measured. The area under this curve, namely area under receiver operator characteristic curve (AUROC), is calculated for each uncertainty/confidence metric. The AUROC score equals the probability that a randomly chosen incorrect response has a higher uncertainty than a randomly chosen correct response (and vice versa for confidence). A perfect uncertainty/confidence metric would have an AUROC score of 1 while a random metric would have 0.5. As an additional performance metric, the area under the Precision-Recall curve (AUPR) is also calculated. The average AUPR scores over two setups, i.e., using either correct or incorrect samples as the positive class, are reported. Note that a higher semantic density corresponds to a higher confidence.
@@
The performance of semantic density (SD) was compared with six existing LLM uncertainty quantification methods: semantic entropy (SE) [23], P(True) [22], degree (Deg) [27], length-normalized likelihood (NL) [32], length-normalized entropy (NE) [30] and predictive entropy (PE) [22]. These methods were applied to seven SOTA open-source LLMs: Llama-2-13B [46], Llama-2-70B [46], Llama-3-8B [3], Llama-3-70B [3], Mistral-7B [19], Mixtral-8x7B [20] and Mixtral-8x22B [44]. Each LLM was tested on four free-form question-answering datasets commonly used in the literature: CoQA [39], TriviaQA [21], SciQ [49] and Natural Questions (NQ) [24]. For each question, 10 responses were generated using group beam search and used as reference responses in calculating SD, SE, Deg, NE, and PE (note that P(True) and NL do not need reference responses). Each unique response among these 10 will also be used as a target response, i.e., the response that needs an uncertainty/confidence estimation, in calculating the AUROC scores of uncertainty/confidence metrics. Detailed experimental configuration and parametric setup is described in Appendix A.1.
@@
Table 1 and Table A2 (in Appendix A.4) show the AUROC and AUPR scores of each uncertainty/confidence metric across different models and datasets, with the best entry in each configuration highlighted in boldface. SD performs best in 26 out of 28 cases for AUROC and 27 out of 28 cases for AUPR, demonstrating that it is reliable and robust as a confidence metric for LLM responses. For AUROC, in two cases it is outperformed by Deg. After investigation, the inherent sequence likelihood returned by the original LLM was badly calibrated in these two cases. Deg is the only method that ignores the likelihood information during its calculation, making its performance unaffected by this negative factor. However, for the other 26 cases, SD is able to utilize the likelihood information to its advantage and outperform Deg. Appendix A.5 includes the results of another variant of SE [11], which is comparable to the original version. SD outperforms it in all the test cases.
@@
To confirm that the observed performance differences in Table 1 are statistically significant, a paired $t$-test (paired by LLM and dataset) was performed between SD and the other metrics (Table A1 in Appendix A.2). The $p$-values are consistently below $10^{-6}$, indicating that the performance gains of SD are strongly statistically significant.
The experiments reported above use the same setup as Kuhn et al. [23] to evaluate the correctness of responses: If the Rouge-L [25] between one response and the reference answer is larger than 0.3, then the response is deemed as correct. In order to investigate the influence of the correctness-checking criterion, the threshold of Rouge-L was varied between 0.1 and 1.0; the resulting performance of different uncertainty/confidence metrics are shown in Figure A1. SD consistently outperforms other methods under the different Rouge-L thresholds. Moreover, when the Rouge-L threshold increases, the AUROC scores generally increase for SD whereas they decrease for most other methods. Since a higher Rouge-L threshold means stricter correctness checking, the experimental performance gain of SD may be even larger if such checking is further improved (e.g. by using a more capable LLM).
To further evaluate the generalizability of SD in other types of tasks, an empirical study using a summarization task (DUC-2004 [36]) was performed. Appendix A.6 shows the performance comparisons among different uncertainty/confidence metrics. SD performs best in most test cases, demonstrating good generalizability.
@@ 4.2 Robustness of Semantic Density
In real-world applications, users may have different preferences when generating responses: Some may prefer a greedy sampling strategy while others may need diverse responses. The second study thus investigated how each uncertainty/confidence metric performs when the target response is sampled using different such strategies. The diverse beam search method inherently utilizes different strategies for each beam group: The first group performs a greedy beam search while later groups encourage more diverse responses. Following Section 4.1, the AUROC scores were calculated for target responses from each group separately, and the results averaged over the four datasets.
@@
Figure 2: AUROC scores over the different beam groups. The plots show the average AUROC scores over the four datasets for the different beam groups in diverse beam search. A smaller group index corresponds to the group with a more greedy generation strategy, while the group with a larger index tends to be more diverse during response generation. Each subfigure corresponds to one of the seven LLMs, as indicated by the subfigure heading. Each curve represents one uncertainty/confidence metric indicated by the legend at the lower right. Semantic density exhibits consistently better and stable performance across different groups, compared to other methods.
@@ 5 Discussion and Future work
One limitation of semantic density is that it needs access to the output probabilities of generated tokens, which may not be available in some proprietary LLMs. In such a case, the more expensive variant in Eq. 5 can be considered as an alternative. Since semantic density does not require any further access to the internal states or weights of the original LLMs, it is still widely applicable. Another limitation is that most responses in the current experiments are at the sentence level. However, an extension of semantic density to long-paragraph responses should be feasible. Following the existing solutions [29, 11], a long response can be decomposed into sentence-level claims [29] or factoids [11], and semantic density then applied to estimate the confidence of each claim/factoid.
The framework for measuring semantic density is modular, and therefore its main components can be extended in future work. First, new sampling strategies that explicitly encourage a better coverage of semantic space can be developed to generate reference responses. Such extensions should improve the reliability of semantic density further. Second, text embedding methods that can measure contextual semantic similarity between responses more reliably will be helpful as well. Third, kernel functions specifically designed for the semantic space should allow the utilization of semantic relationships more efficiently. Fourth, more precise methods for calibrating inherent token probabilities will form a more reliable base for calculating semantic density.
@@ 6 Conclusion
This paper proposes semantic density as a practical new metric for measuring the confidence of LLM responses. It overcomes the limitations of existing approaches by utilizing the semantic information in an efficient and precise manner. It is response-specific, "off-the-shelf", and applicable to free-form generation tasks. Experimental comparisons with six existing uncertainty/confidence metrics across seven SOTA LLMs and four benchmark datasets suggest that it is accurate, robust, and general, and can therefore help deploy LLMs in safety-critical domains.
@@ A.1 Experimental Setup
This section provides the details of the experimental setup for reproducing the results presented in Section 4. The source codes for reproducing the reported experimental results are provided at: https://github.com/cognizant-ai-labs/semantic-density-paper.
@@
Correctness Checking:
Following the setup in Kuhn et al. [23], for all the reported experimental results except Figure A1, a response is considered to be correct if the Rouge-L [25] between it and any of the reference answers is larger than 0.3, after trimming the redundant continuations.
@@
The details of the datasets are as follows:
@@
[…]
Uncertainty/Confidence Metrics:
The same exact target response and reference responses are used for all the tested methods. For SD, SE, and Deg, the microsoft/deberta-large-mnli model from Huggingface Transformers library [50] is used as the NLI classification model. Following Kuhn et al. [23], the probabilities for "contradiction", "neutral" and "entailment" are averaged bidirectionally. The parametric setup for the uncertainty metrics tested in Section 4 is summarized as follows:
  • semantic density (SD): The same exact steps as described in Algorithm 1 were implemented, with a fixed temperature of 0.1 applied to rescale the value of each token probability during postprocessing.
@@
  • degree (Deg): The "entailment" probability returned by the NLI model (averaged bidirectionally) is used as the similarity between two responses, which is then used as the diagonal element in the degree matrix.
@@ A.2 Results of Statistical Tests
Table A1 shows the $p$-values of the paired $t$-tests as described in Section 4.1.
Table A1: Statistical significance of SD’s advantage over other methods ($p$-values of paired $t$-tests)
@@
A.3 Performances with Different Correctness Thresholds
Figure A1 shows the performance changes of the uncertainty/confidence metrics with different Rouge-L thresholds when checking the correctness of responses. SD performs the best overall across these different criteria.
Figure A1: Performance Evaluation over Different Correctness Checking Criteria. The Rouge-L threshold for checking response correctness is set as 0.1, 0.3, 0.5, 0.7, 0.9 and 1.0, respectively. Each subfigure corresponds to one combination of base LLM and dataset, indicated by the subfigure heading. Each curve corresponds to the performance of one uncertainty/confidence metric identified in the legend at the very top and bottom. For each subfigure, the horizontal axis indicates the Rouge-L threshold, and the vertical axis indicates the AUROC score. Overall, semantic density exhibits best performance under different Rouge-L thresholds. The experimental performance of SD increases when the correctness-checking criterion becomes stricter (a higher Rouge-L threshold), while most other methods show opposite trends.
A.4 Performance Evaluation Using AUPR
During this evaluation, each uncertainty/confidence metric was used as a quantitative detector in two cases: 1) detecting incorrect responses, 2) detecting correct responses. The AUPR scores for both cases were calculated. Table A2 shows the average AUPR scores for them. Again, SD performs the best overall.
Table A2: Performance evaluation using AUPR-average metric
CoQA
AUPR SD SE P(True) Deg NL NE PE
Llama-2-13B 0.766 0.605 0.589 0.724 0.692 0.617 0.632
Llama-2-70B 0.766 0.587 0.566 0.710 0.686 0.602 0.628
Llama-3-8B 0.718 0.581 0.579 0.779 0.665 0.598 0.595
Llama-3-70B 0.765 0.572 0.641 0.714 0.658 0.566 0.606
Mistral-7B 0.773 0.600 0.652 0.730 0.680 0.599 0.618
Mixtral-8x7B 0.768 0.592 0.580 0.719 0.686 0.599 0.636
Mixtral-8x22B 0.772 0.581 0.605 0.715 0.672 0.585 0.632
TriviaQA
AUPR SD SE P(True) Deg NL NE PE
Llama-2-13B 0.839 0.651 0.589 0.807 0.662 0.570 0.550
Llama-2-70B 0.822 0.637 0.543 0.775 0.670 0.572 0.552
Llama-3-8B 0.855 0.636 0.636 0.781 0.815 0.638 0.616
Llama-3-70B 0.817 0.609 0.629 0.750 0.769 0.594 0.573
Mistral-7B 0.857 0.668 0.586 0.813 0.720 0.607 0.529
Mixtral-8x7B 0.835 0.653 0.546 0.782 0.763 0.628 0.591
Mixtral-8x22B 0.815 0.630 0.573 0.747 0.740 0.616 0.586
SciQ
AUPR SD SE P(True) Deg NL NE PE
Llama-2-13B 0.744 0.567 0.559 0.695 0.686 0.511 0.556
Llama-2-70B 0.733 0.626 0.574 0.694 0.630 0.549 0.604
Llama-3-8B 0.768 0.605 0.560 0.714 0.684 0.590 0.646
Llama-3-70B 0.755 0.597 0.551 0.690 0.722 0.548 0.518
Mistral-7B 0.759 0.612 0.559 0.717 0.664 0.561 0.530
Mixtral-8x7B 0.759 0.612 0.570 0.702 0.726 0.605 0.645
Mixtral-8x22B 0.753 0.604 0.588 0.703 0.700 0.589 0.628
NQ
AUPR SD SE P(True) Deg NL NE PE
Llama-2-13B 0.656 0.561 0.552 0.642 0.589 0.544 0.598
Llama-2-70B 0.655 0.539 0.525 0.648 0.580 0.547 0.591
Llama-3-8B 0.663 0.555 0.512 0.644 0.596 0.563 0.572
Llama-3-70B 0.700 0.562 0.608 0.655 0.638 0.570 0.580
Mistral-7B 0.654 0.573 0.519 0.637 0.623 0.586 0.586
Mixtral-8x7B 0.700 0.590 0.555 0.692 0.650 0.576 0.583
Mixtral-8x22B 0.687 0.571 0.509 0.671 0.645 0.596 0.647
A.5 Performance of Another Variant of Semantic Entropy
Table A3 shows the performance of a follow-up variant of semantic entropy [11]. The performance of this variant is comparable with the original semantic entropy, with several cases worse than the original version. The proposed semantic density outperforms this variant in all test cases.
Table A3: Performance of the semantic entropy variant
AUROC CoQA TriviaQA SciQ NQ
Llama-2-13B 0.587 0.726 0.576 0.576
Llama-2-70B 0.570 0.696 0.596 0.555
Mistral-7B 0.606 0.736 0.587 0.582
Meta-Llama-3-8B 0.571 0.694 0.582 0.563
Meta-Llama-3-70B 0.572 0.734 0.613 0.568
Mixtral-8x7B 0.567 0.699 0.576 0.612
Mixtral-8x22B 0.562 0.690 0.592 0.565
A.6 Performace Evaluation in a Summarization Task
This experiment used Task 1 of DUC-2004 dataset [36], which is a single-document summarization task. The LLMs are prompted to summarize the document using one sentence. Table A4 presents the results. SD performs the best in six out of seven test cases, demonstrating its good generalizability to tasks other than question-answering.
Table A4: Performance of different uncertainty metrics for summarization task
DUC 2004
AUROC SD SE Deg NL NE PE
Llama-2-13B 0.622 0.612 0.606 0.594 0.601 0.521
Llama-2-70B 0.552 0.579 0.549 0.520 0.575 0.534
Llama-3-8B 0.597 0.529 0.590 0.538 0.538 0.521
Llama-3-70B 0.593 0.513 0.592 0.489 0.515 0.433
Mistral-7B 0.679 0.537 0.660 0.673 0.551 0.549
Mixtral-8x7B 0.640 0.532 0.620 0.636 0.576 0.497
Mixtral-8x22B 0.649 0.625 0.639 0.574 0.630 0.532